OpenStudy (ninja123987):

what is the square root of 36

OpenStudy (lyss):

6*6 = 36

HanAkoSolo (jamierox4ev3r):

\(\Large\sqrt{36}\) \(\Large 6*6=6^2\)

OpenStudy (3mar):

\[\Huge\color{purple} {\sqrt{(36)}=\pm6}\]

KjSaif (kjsaif):

since 6 times 6=36, the square root of 36 is 6

HanAkoSolo (jamierox4ev3r):

@3mar bring a good point to the table \(\Large (-6)^2 =36\) \(\Large (6)^2 =36\)

OpenStudy (3mar):

Thanks, @Jamierox4ev3r This is the definition of the square root! What is the numbers that if they squared would give the number under the square root?? \(-6~~and ~~-6\) give 36 if they are squared!

OpenStudy (adelyn_helps):

\[6\times6=36 \sqrt{36}=6\]